OpenStudy (anonymous):

write the point-slope form of the line described: through (0,-1) and parallel to y=3/4x-4

OpenStudy (anonymous):

The slope is 3/4 so what you do Is solve for y using this formula.\[y +1 = \frac{ 3 }{ 4 }(x)\]

OpenStudy (goformit100):

Why am I tagged here , I don't know the answer, Sorry Madam.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@mebs how do i solve for y?

OpenStudy (anonymous):

subtract 1 from both sides to get \[y=\frac{3}{4}x-1\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then what? @satellite73

OpenStudy (anonymous):

parallel lines will have the same slope. So we know the slope(m) is 3/4. point slope formula : y - y1 = m(x - x1) slope(m) = 3/4 (0,-1) x1 = 0, and y1 = -1 now sub that into the formula and you will have your answer :)

OpenStudy (mandre):

The point slope form is: \[y - y _{1}= m(x - x _{1})\] Using point (0,-1) and parallel line y=3/4x-4 you have: \[y-(-1) = \frac{ 3 }{ 4 }(x - 0)\] Since they ask for the point-slope form that would be your answer or the simplified form: \[y + 1 = \frac{ 3 }{ 4 }x\]